# Holographic entanglement negativity and thermodynamics in backreacted AdS black hole

### v1

@@ 9 Summary and Discussion
[…]
Overall, the observations made here provide a deeper understanding of entanglement and related phenomena in strongly coupled conformal field theories with backreaction at finite temperature. These results establish that the influence of heavy flavour backreaction is not universal but strongly regime dependent as it reduces entanglement in the vacuum like phase and enhances it in the thermal phase. This dual role highlights HEN as a sharper diagnostic of the balance between quantum and thermal correlations than HEE alone. In our study of entangled-disentangled phase transition using HEN for disjoint configuration we found that the critical separation decreases with backreaction. Also the critical separation from HEN ($l_{c}^{HEN}$) is always smaller then obtained from HMI ($l_{c}^{HMI}$) i.e. $l_{c}^{HEN}<l_{c}^{HMI}$.
Future directions include performing a systematic numerical study in the intermediate regime $T_{f}l\sim\mathcal{O}(1)$ to map out crossover behaviour, exploring time-dependent quenches to connect with thermalization dynamics and extending the analysis to higher-derivative gravity or string-theoretic embeddings to quantify $1/N$ and $\alpha^{\prime}$ corrections. Such studies could shed further light on the microscopic structure of entanglement in strongly coupled plasmas with flavour backreaction and on the universal features of entanglement negativity in holographic duals.
@@ Appendix A Numerical coefficients
The constants used in eq.˜4.46 are given by

### v2

@@ 9 Summary and Discussion
To summarize, we have investigated the entanglement thermodynamics and holographic entanglement negativity (HEN) in a strongly coupled large $N_{c}$ gauge theory at finite temperature backreacted by the heavy static fundamental quarks. This field theory is holographically dual to a deformed AdS black hole geometry where the deformation/backreaction is sourced by the large number of long static strings (string cloud). We began by revisiting the holographic entanglement entropy in both the low and high temperature regimes. The analysis of HEE in low temperature limit allowed us to identify a first law like relation for the entanglement entropy. Furthermore, we provided a review of the entanglement wedge cross section and discussed its implications for reflected entropy.
Subsequently, we explored the behavior of HEN for adjacent, bipartite and disjoint subsystem configurations. The main focus of this work is to understand how the backreaction modifies quantum correlations across different thermal regimes and how these results compare with the corresponding behavior of HEE. Our analysis reveals a striking regime dependence: for adjacent subsystems case, the low temperature expansion produces a correction proportional to $(1-\rho)T_{f}^{\,d}$ which enhances HEN and thus confirms that backreaction increases the distillable entanglement in the near vacuum phase. In the high temperature limit, the relevant $\widetilde{S}(\rho,d)$ coefficients increase with $\rho$ leading to an enhancement of HEN. For the bipartite configuration, the HEN shows increasing behavior with backreaction in low and high temperature limits, and isolate the thermal entropy. For disjoint subsystems, a similar increasing trend emerges in the low $T_{f}$ regime is dominated by $(1-\rho)$ terms, while in the high $T_{f}$ due to monotonically increasing $\widetilde{S}(\rho,d)$ coefficients. In contrast to the holographic entanglement entropy and mutual information which capture total correlations and become insensitive to quantum entanglement at high temperature, the more reliable measure provided by holographic entanglement negativity successfully captures the enhancement of distillable quantum correlations induced by backreaction. From our analysis of the critical separation for the disjoint configuration using holographic entanglement negativity, we find that the critical separation increases with increasing backreaction. This behavior is consistent with the enhancement of quantum correlations induced by backreaction, implying that a larger spatial separation is required to completely destroy entanglement. In this sense, stronger backreaction shifts the transition between the entangled and disentangled phases to larger separations, highlighting the role of matter backreaction in shaping the entanglement structure of strongly coupled holographic plasmas.
Overall, the observations made here provide a deeper understanding of entanglement and related phenomena in strongly coupled conformal field theories with backreaction at finite temperature. These results establish that the influence of adding the heavy flavor backreaction is universal and increases the entanglement in the theory through the extra degrees of freedom. This highlights HEN as a sharper diagnostic of the balance between quantum and thermal correlations than HEE alone. Moreover, our findings are consistent with the previous studies on the effect of backreaction on entanglement. Future directions include performing a systematic numerical study in the intermediate regime $T_{f}l\sim\mathcal{O}(1)$ to map out crossover behavior, exploring time-dependent quenches to connect with thermalization dynamics and extending the analysis to higher-derivative gravity or string-theoretic embeddings to quantify $1/N$ and $\alpha^{\prime}$ corrections. Such studies could shed further light on the microscopic structure of entanglement in strongly coupled plasmas with flavor backreaction and on the universal features of entanglement negativity in holographic duals.
@@ Appendix A Numerical coefficients
The constants used in eq.˜4.47 are given by